# Predictively Oriented Posteriors

### v1

@@ Acknowledgements
This manuscript has benefited tremendously from interactions with many brilliant colleagues. Most notably, we feel deep gratitude towards Prof. Chris Oates for many conversations and insightful remarks. Further, we would like to thank Dr. Louis Sharrock and Dr. Sam Power for their help with the computational aspects of this paper. We are also grateful for correspondence with Prof. Pierre Alquier that pointed to the connection with PAC-Bayesian majority voting. Finally, we want to thank Dr. Evgenii Chzhen for insightful discussions on sub-Gaussianity. DT Frazier acknowledges funding from the Australian Research Council (DE200101070, DP200101414).
@@ Appendix A Notations
In this section we re-describe the main notations introduced in the main text to ensure the appendix is as self-contained as possible. We consider observed data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$, where each $x_{i}$ is generated from an unknown distribution $P_{0}$. We denote expectations of functions $g:\mathcal{X}\to\mathbb{R}$ with respect to $X\sim P_{0}$ by $\mathbb{E}[g(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[g(X)]$ if we wish to emphasise the distribution over which integration occurs. The model class is given by $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior measure $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$. Likewise, let $\mathcal{P}(\Theta)$ denote the set of distributions over $\Theta$, and recall $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$ for $Q\in\mathcal{P}(\Theta)$. The function $S(P,x)$ denotes a scoring rule for the distribution $P$, while $\mathcal{S}(P,P)=\mathbb{E}_{X\sim P}[S(P,X)]$ denotes the expected score, and $\mathcal{D}_{S}(P,P_{0})=\mathcal{S}(P,P_{0})-\mathcal{S}(P_{0},P_{0})$ is the divergence associated to $P$, evaluated at the true distribution $P_{0}$. Further, define $S_{n}(P_{\theta})=\frac{1}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})$, and $L_{n}(\theta_{1:k})=\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$, where $L(\theta_{1:k},x)$ is as defined in Assumption 2, and which differs by the choice of approximation.
@@ Appendix B Lemmas
To prove this, take $x,x^{\prime}\in\mathcal{X}$, let $\kappa(x,x^{\prime})$ denote a positive-definite kernel function with feature map $\phi:\mathcal{X}\to\mathcal{H}$, such that $\kappa(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}}$. For fixed $\theta\in\Theta$, define the mean embedding $\mu_{\theta}\in\mathcal{H}$ by $\mu_{\theta}:=\int\phi(z)\,\mathsf{d}P_{\theta}(z)$ and the inner product $\langle\mu_{\theta},\mu_{\theta^{\prime}}\rangle_{\mathcal{H}}=\int\!\!\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime}).$ Following Gneiting et al. (2007), kernel scoring rules can be expressed as $S(P,x)=\mathbb{E}_{X,X^{\prime}\sim P}[\kappa(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[\kappa(X,x)]+\kappa(x,x).$
@@
The map $P_{\theta}\mapsto\mu_{\theta}$ is linear, so that $P_{\theta}\mapsto\|\mu_{\theta}\|_{\mathcal{H}}^{2}$ is quadratic. Thus, $P_{\theta}\mapsto S(P_{\theta},x)$ is the composition of a linear map and a convex function, and is convex. To deduce the second result, write
@@
The log score satisfies Assumption 1. Further, for all $x\in\mathcal{X}$,
                                                                                                    $-\log\int p_{\theta}(x)\mathsf{d}Q(\theta)$ ${}={}$ $-\int\log p_{\theta}(x)\mathsf{d}Q(\theta)-\Delta_{\log}(Q,x)$                                                                                                      (29)
where for $g(p_{\theta}(x),p_{Q}(x))$ some intermediate value between $p_{\theta}(x)$ and $\int p_{\theta}(x)\mathsf{d}Q(\theta)$, we have
                                                                                                    $\Delta_{\log}(Q,x)$ ${}={}$ $\int\frac{1}{2g(p_{\theta}(x),p_{Q}(x))^{2}}\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}\mathsf{d}Q(\theta).$                                                                                                      (30)
Proof of Lemma 4.
For $\mu(x)=\int p_{\theta}(x)\mathsf{d}Q(\theta)$, the function $\log\mu(x)$ is convex, since $\log\mu(x)$ is concave. Thus, Assumption 1 is satisfied for the log score. Further, it is well-known that the KL-divergence is jointly convex in both arguments.
To prove the second part of the result, let $y(x)=p_{\theta}(x)$, and consider a second-order Taylor expansion of $\log y(x)$, around $\mu(x)$: for $g(y(x),\mu(x))$ is a real value between $y(x)$ and $\mu(x)$,
$\log y(x)=$ $\log\mu(x)+\frac{1}{\mu(x)}(y(x)-\mu(x))-\frac{1}{2g(y(x),\mu(x))^{2}}(y(x)-\mu(x))^{2}$
$=$ $\log\mathbb{E}_{\theta^{\prime}\sim Q(\theta^{\prime})}[p_{\theta^{\prime}}(x)]-\frac{1}{2g\{p_{\theta}(x),\int p_{\theta}(x)\mathsf{d}Q(\theta)\}^{2}}\left\{p_{\theta}(x)-\mathbb{E}_{\theta^{\prime}\sim Q(\theta^{\prime})}[p_{\theta^{\prime}}(x)]\right\}^{2}.$
Therefore,
$-\log\int p_{\theta}(x)\mathsf{d}Q(\theta)=$ $-\int\log p_{\theta}(x)\mathsf{d}Q(\theta)-\int\frac{\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}}{2g\{p_{\theta}(x),p_{Q}(x)\}^{2}}\mathsf{d}Q(\theta)$
$\Delta_{\log}(Q,x)=$ $\int\frac{\left\{p_{\theta}(x)-\int p_{\theta^{\prime}}(x)\mathsf{d}Q(\theta^{\prime})\right\}^{2}}{2g\{p_{\theta}(x),p_{Q}(x)\}^{2}}\mathsf{d}Q(\theta).$
Lemma 5.
Under Assumptions 1-4, for any $\lambda_{n}\geq d_{S}/r_{0}$, we have
@@
Furthermore, for the specific choice $\lambda_{n}=n^{1/2}\,\sqrt{\frac{d_{S}}{C^{2}}\,\log\left(\frac{ec}{d_{S}}\right)}$, for any $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, we have
@@
The proof follows similar arguments to those used to prove Lemma 1 but where we take $L(\theta,\theta^{\prime})=R(\theta)=\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})$ and $L_{n}(\theta,\theta^{\prime})=R_{n}(\theta)=S_{n}(P_{\theta})-S_{n}(P_{0}).$ If Assumption 3 is satisfied for $R(\theta)$ and $R_{n}(\theta)$, then
@@
$=$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}$ (32)
@@
$\pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\pi(\theta)}{\Pi(\mathcal{B}_{r})}\quad\textnormal{with}\quad\mathcal{B}_{r}:=\left\{\theta\in\Theta:\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}.$
@@ Appendix C Proofs of Main results
Let us consider any fixed integer $k>1$. Recall $Q^{k}=\bigotimes_{j=1}^{k}Q$. Apply the Donsker-Varadhan Lemma with $h(\theta_{1:k})=k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}$, to see that, $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$, we have
@@
The last term on the right-hand-side is controlled using Tonelli’s theorem and Assumption 3, under which:
$\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})=\log\mathop{\text{\large$\int$}}\nolimits\mathbb{E}e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})\leq\frac{k^{2}\lambda_{n}^{2}C_{k}^{2}}{n},$
@@
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
@@
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
To obtain the first result of the theorem, we just need to lower bound the expected integral on the left-hand-side. From the convexity in Assumption 1, for the DI-based approximation (Assumption 2 (DI)), we have
@@
where $\delta(\theta_{1:k})=\mathbb{E}[\delta(\theta_{1:k},X)]$; alternatively, by Assumption 1, under the MS-class of approximations (Assumption 2 (MS)), we have
$\mathcal{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)$ $\leq\int\mathcal{S}\left(\frac{1}{k}\sum_{j=1}^{k}P_{\theta_{j}},P_{0}\right)\mathsf{d}Q_{n}^{k}(\theta_{1:k})\equiv\int L(\theta_{1:k})\mathsf{d}Q_{n}(\theta_{1:k}).$
Finally, and recalling $P_{Q_{n}}=\int P_{\theta}\mathsf{d}Q_{n}(\theta)$, this gives us
$\mathbb{E}\left[\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$ (33)
as desired.
To obtain the second result, notice that when no approximation is used, we have for any $Q$ and any $x$ that
$S\left(P_{Q},x\right)=\widehat{S}\left(P_{Q},x\right)\equiv\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),$
which implies that for any $Q$, we have
$S_{n}(P_{Q})\equiv\frac{1}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)=\frac{1}{n}\sum_{i=1}^{n}\int L(\theta_{1:k},x_{i})\mathsf{d}Q^{k}(\theta_{1:k})\equiv\int L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k}).$
Using the above, we can rewrite equation (33) as
$\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$
Recalling that $\mathcal{S}(P_{0},P_{0})=\mathbb{E}S_{n}(P_{0})$, we can subtract $\mathcal{S}(P_{0},P_{0})$ from both sides and write
@@
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})-\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$

### v2

@@ Acknowledgements
This manuscript has benefited tremendously from interactions with many brilliant colleagues. Most notably, we feel deep gratitude towards Prof. Chris Oates for many conversations and insightful remarks. Further, we would like to thank Dr. Louis Sharrock and Dr. Sam Power for their help with the computational aspects of this paper. We are also grateful for correspondence with Prof. Pierre Alquier that pointed to the connection with PAC-Bayesian majority voting. Additionally, we owe a tremendous debt of gratitude to several colleagues who took time to read an earlier version of this manuscript and gave incredibly insightful comments—often pages and pages of them!—that significantly improved the paper; most notably Prof. David Dunson, Prof. Jeff Miller, Prof. Aki Vehtari, Prof. Pierre Jacob, Dr. Andres Masegosa, Dr. Edwin Fong, Dr. Jack Jewson, and Dr. Takuo Matsubara. DTF acknowledges funding from the Australian Research Council (DE200101070, DP200101414), YM from the Heilbronn Institute for Mathematical Research (EP/V521917/1), and JK was supported through the UK’s Engineering and Physical Sciences Research Council (EPSRC) via EP/W005859/1. The authors acknowledge the use of the UCL Myriad High Performance Computing Facility (Myriad@UCL), and associated support services, in the completion of this work. Data were provided by the UK National River Flow Archive.
@@ Appendix A Notations
In this section we re-describe the main notations introduced in the main text to ensure the appendix is as self-contained as possible. We consider observed data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$, where each $x_{i}$ is generated from the true unknown distribution $P_{0}$. We denote expectations of functions $g:\mathcal{X}\to\mathbb{R}$ with respect to $X\sim P_{0}$ by $\mathbb{E}[g(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[g(X)]$ if we wish to emphasise the distribution over which integration occurs. The model class is given by $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior measure $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$. Likewise, let $\mathcal{P}(\Theta)$ denote the set of distributions over $\Theta$, and recall $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$ for $Q\in\mathcal{P}(\Theta)$. The function $S(P,x)$ denotes a scoring rule for the distribution $P$, while $\mathcal{S}(P,P)=\mathbb{E}_{X\sim P}[S(P,X)]$ denotes the expected score, and $\mathcal{D}_{S}(P,P_{0})=\mathcal{S}(P,P_{0})-\mathcal{S}(P_{0},P_{0})$ is the divergence associated to $P$, evaluated at the true distribution $P_{0}$. Further, define $S_{n}(P_{\theta})=\frac{1}{n}\sum_{i=1}^{n}S(P_{\theta},x_{i})$, and $L_{n}(\theta_{1:k})=\frac{1}{n}\sum_{i=1}^{n}L(\theta_{1:k},x_{i})$, where $L(\theta_{1:k},x)$ is as defined in Assumption 2.
@@ Appendix B Lemmas
To prove this, take $x,x^{\prime}\in\mathcal{X}$, let $\kappa(x,x^{\prime})$ denote a positive-definite kernel function with feature map $\phi:\mathcal{X}\to\mathcal{H}$, such that $\kappa(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}}$. For fixed $\theta\in\Theta$, define the mean embedding $\mu_{\theta}\in\mathcal{H}$ by $\mu_{\theta}:=\int\phi(z)\,\mathsf{d}P_{\theta}(z)$ and the inner product $\langle\mu_{\theta},\mu_{\theta^{\prime}}\rangle_{\mathcal{H}}=\int\!\!\int\kappa(z,z^{\prime})\,\mathsf{d}P_{\theta}(z)\mathsf{d}P_{\theta^{\prime}}(z^{\prime}).$ Following Gneiting et al. (2007), all kernel scoring rules can be expressed as $S(P,x)=\mathbb{E}_{X,X^{\prime}\sim P}[\kappa(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[\kappa(X,x)]+\kappa(x,x).$
@@
The map $P_{\theta}\mapsto\mu_{\theta}$ is linear, so that $P_{\theta}\mapsto\|\mu_{\theta}\|_{\mathcal{H}}^{2}$ is quadratic. Thus, $P_{\theta}\mapsto S(P_{\theta},x)$ is the composition of a linear map and a convex function, and so is convex. To deduce the second result, write
@@
Under Assumptions 1-3, for any $\lambda_{n}\geq d_{S}/r_{0}$, we have
@@
Furthermore, for the specific choice $\lambda_{n}=n^{1/2}\,\sqrt{{d_{S}}/{C^{2}}\,\log\left({e\cdot c_{S}}/{d_{S}}\right)}$, for any $n\,\geq\frac{d_{S}C^{2}}{r_{0}^{2}\,\log\left(\frac{e\cdot c_{S}}{d_{S}}\right)}$, we have
@@
The proof follows similar arguments to those used to prove Lemma 1 but where we take $L(\theta,\theta^{\prime})=R(\theta)=\mathcal{S}(P_{\theta},P_{0})-\mathcal{S}(P_{0},P_{0})$ and $L_{n}(\theta,\theta^{\prime})=R_{n}(\theta)=S_{n}(P_{\theta})-S_{n}(P_{0}).$ If Assumption 2 is satisfied for $R(\theta)-R_{n}(\theta)$, then
@@
$=$ $\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\left\{S_{n}(P_{\theta})-S_{n}(P_{0})\right\}\mathsf{d}Q(\theta)+\frac{\mathrm{KL}({Q\|\Pi})}{\lambda_{n}}\right\}+\frac{\lambda_{n}C^{2}}{n}.$ (21)
@@
$\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r})}\quad\textnormal{with}\quad\mathcal{B}_{r}:=\left\{\theta\in\Theta:\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}.$
@@
B.1 Existence and uniqueness of the minimiser
It is interesting to investigate the sufficient conditions under which the objective of (6) admits a unique minimiser. We posit two assumptions in order to do so: one on the score and the other on the prior.
Lemma 4.
Suppose that Assumption 1 holds and further that the score is bounded from below so that $\inf_{Q\in\mathcal{P}(\Theta)}\mathscr{L}(Q)=M>$-$\infty$. Assume further that the prior $\Pi$ admits density with respect to the Lebesgue measure $\mathsf{d}\Pi(\theta)\propto\exp\{-U(\theta)\}$ for $U:\Theta\to\mathbb{R}$ such that:
  1. (i)
    there exist constants $c>0$ and $c^{\prime}\in\mathbb{R}$ such that for all $\theta\in\Theta$,
    $\nabla_{\theta}U(\theta)\cdot\theta\geq c\lVert\theta\rVert^{2}+c^{\prime},$
    and,
  2. (ii)
    $\nabla_{\theta}U$ is Lipschitz continuous.
Then there exists a unique measure $Q_{n}\in\mathcal{P}(\Theta)$ such that $Q_{n}=\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathscr{L}$, and further this $Q_{n}$ is characterised by the implicit equation
$\mathsf{d}Q_{n}(\theta)\propto\mathsf{d}\Pi(\theta)\cdot\exp\left\{\left.-\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)\right\}.$
Proof.
We have by Assumption 1 that the score $S(\cdot,\,x)$ is convex is its first argument. In combination with the fact that the mapping $Q\mapsto P_{Q}$ is linear, we have the the mapping $Q\mapsto S(P_{Q},\,x)$ is convex. The existence and uniqueness of a minimiser is then provided under the same conditions in Proposition 2.5 of Hu et al. (2020), and we point the reader to their Section 4 for a proof.
To characterise the form of this minimiser, we take the functional derivative of (6) with respect to $Q$, set it equal to zero, and solve for $Q$. We begin by noting that the functional derivative of the KL regulariser evaluated at $\theta$ is given by
$1+\log\left\{\frac{\mathsf{d}Q}{\mathsf{d}\Pi}(\theta)\right\},$ (22)
while that of the score term is
$\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\delta S(P_{Q},x_{i})}{\delta Q}(\theta).$ (23)
By the linearity of functional derivatives, setting the sum of (22) and (23) equal to zero and removing constant terms, we find that $Q_{n}$ must satisfy
$0=\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\left.\frac{\delta S(P_{Q},x_{i})}{\delta Q}\right|_{Q=Q_{n}}(\theta)+\log\left\{\frac{\mathsf{d}Q_{n}}{\mathsf{d}\Pi}(\theta)\right\}$
whence rearranging yields the stated result. ∎
The prior coercivity conditions of i and ii are satisfied, for instance, by choosing the prior to be a multivariate standard Gaussian prior, common to many applications (Hu et al., 2020). And the condition that the infimum be attained at non-infinite $M$ is immediate, for example, for the squared MMD under bounded kernels (see also Shen et al., 2025, Theorem 1). In the case of the log score, we refer the reader to Liu et al. (2025, Proposition 1) who establish the necessary conditions.
@@ Appendix C Proofs of Main results
The result depends on slightly different arguments depending on whether Assumption 2 or Assumption 2 is satisfied. Therefore, we break the result into cases.
Scores Satisfying Assumption 2. Consider any fixed integer $k>1$. Recall $Q^{k}=\bigotimes_{j=1}^{k}Q$. Apply the Donsker-Varadhan Lemma with $h(\theta_{1:k})=k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}$, to see that, $P_{0}$-almost surely, $\forall Q\in\mathcal{P}(\Theta)$, we have
@@
The last term on the right-hand-side is controlled using Tonelli’s theorem and Assumption 2(ii), under which:
$\log\mathbb{E}\mathop{\text{\large$\int$}}\nolimits e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})=\log\mathop{\text{\large$\int$}}\nolimits\mathbb{E}e^{k\lambda_{n}\{L(\theta_{1:k})-L_{n}(\theta_{1:k})\}}\mathsf{d}\Pi^{k}(\theta_{1:k})\leq\frac{k^{2}\lambda_{n}^{2}C^{2}}{n},$
@@
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ $\leq\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\mathbb{E}\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}+\frac{\lambda_{n}kC^{2}}{n}$
$=\mathbb{E}\left[\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})+\frac{\text{KL}(Q_{n}^{k}\|\Pi^{k})}{k\lambda_{n}}\right]+\frac{\lambda_{n}kC^{2}}{n}.$
@@
$\mathbb{E}\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}.$
Now, we lower bound $\mathop{\text{\large$\int$}}\nolimits L(\theta_{1:k})\mathsf{d}Q_{n}^{k}(\theta_{1:k})$ on the left-hand-side: using the convexity in Assumption 1 and the structure of the loss in Assumption 2(i), we have
@@
where $\delta(\theta_{1:k})=\mathbb{E}[\delta(\theta_{1:k},X)]$. Recalling that $P_{Q_{n}}=\int P_{\theta}\mathsf{d}Q_{n}(\theta)$, this gives us
$\mathbb{E}\left[\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\mathbb{E}\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$ (24)
Using Assumption 2(i), we can rewrite equation (24) as
$\mathbb{E}\left\{\mathcal{S}\left(P_{Q_{n}},P_{0}\right)\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}Q^{k}(\theta_{1:k})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$
Recalling that $\mathcal{S}(P_{0},P_{0})=\mathbb{E}S_{n}(P_{0})$, we subtract $\mathcal{S}(P_{0},P_{0})$ from both sides and write
@@
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{0},P_{0})-\mathcal{S}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}(P_{Q},P_{0})+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC^{2}}{n}.$
Scores Satisfying Assumption 2. For scores that do not satisfy Assumption 2, the above proof technique breaks down. We can, however, derive a similar bound using the variational definition of $Q_{n}$.
From the definition of $\mathcal{D}(P_{Q},P_{0})$, we have
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $=\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]-\mathcal{S}(P_{0},P_{0})$ (25)
$=\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})-\left\{\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]\right\}$
$\leq\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})+\mathbb{E}\sup_{Q\in\mathcal{P}_{\Pi}(\Theta)}|S_{n}(P_{Q})-\mathcal{S}(P_{Q},P_{0})|$
$\leq\mathbb{E}[{S}_{n}(P_{Q_{n}})]-\mathcal{S}(P_{0},P_{0})+r_{n},$ (26)
where the last line follows from Assumption 2. Using the variational definition of $Q_{n}$, we have
$S_{n}(P_{Q_{n}})\leq S_{n}(P_{Q_{n}})+\frac{\mathrm{KL}(Q_{n}\|\Pi)}{\lambda_{n}}=\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}.$
Taking expectations then yields
$\mathbb{E}\left\{S_{n}(P_{Q_{n}})\right\}$ $\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{S_{n}(P_{Q})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}.$ (27)
Plugging equation (27) into (26) yields
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $=\mathbb{E}[\mathcal{S}(P_{Q_{n}},P_{0})]-\mathcal{S}(P_{0},P_{0})$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}-\mathcal{S}(P_{0},P_{0})+r_{n}$
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+r_{n}.$
The stated result now follows from the definition of $\nu_{n}$. ∎